# Drawback of Enforcing Equivariance and its Compensation via the Lens of Expressive Power

### v1

@@ 4.2 LENs further implies Symmetric Channel Vectors
We first prove that each row of $\psi_{g}$ has at most one non-zero entry for all $g\in G$. Assume for contradiction that exists $g\in G$ where a row $i$ of $\psi_{g}$ has at least two non-zero entries, $(\psi_{g})_{ij}$ and $(\psi_{g})_{ik}$. Without loss of generality, assume $(\psi_{g})_{ij}\leq(\psi_{g})_{ik}$. Input vector $x=e_{j}-e_{k}$ gets $(\psi_{g}\sigma(x))_{i}=(\psi_{g}e_{j})_{i}=(\psi_{g})_{ij}$ while $(\sigma(\psi_{g}x))_{i}=\sigma({(\psi_{g})_{ij}-(\psi_{g})_{ik}})=0$. Admitted $\psi$ implies $(\psi_{g})_{ij}=0$, which contradicts our initial assumption that the entries were non-zero. Therefore, each row of $\psi_{g}$ can contain at most one non-zero entry.
@@
Consider for a pair $(i,j)$ making $\psi_{ij}\neq 0$ for some $g\neq h$. Input vector $e_{j}$ gets $\beta_{j}=W^{(2)}e_{j}=W^{(2)}\psi e_{j}=\psi_{ij}W^{(2)}e_{i}=\psi_{ij}\beta_{i}$ for $\psi=\psi_{g}$ and $\psi=\psi_{h}$. Considering the direction of both sides, it shows that $\beta_{i}$ and $\beta_{j}$ has the same direction, meaning $\|\beta_{j}\|\beta_{i}=\|\beta_{i}\|\beta_{j}$. Considering the magnitude, the $\psi_{ij}$ does not change if $\psi$ switches from $\psi_{g}$ to $\psi_{h}$, meaning $(\psi_{g})_{ij}=(\psi_{h})_{ij}$. ∎
In light of this geometric alignment, we define that $\beta_{i}$ and $\beta_{j}$ are in the same orbit if they have the same direction, i.e., $\|\beta_{j}\|\beta_{i}=\|\beta_{i}\|\beta_{j}$. The lemma demonstrates that the invariance constraint forces all weight vectors within the same orbit to share the same direction in vector space. This imposes a strong, predictable structure on the weight space of the network’s final layer, partitioning it into sets of aligned vectors.
Aside from the constraint on $W^{(2)}$, LENs also impose a condition on $W^{(1)}$, which should belong to the intertwiner space defined by $\psi$ and $\rho$. Combined with the constraints on $\psi$ and $W^{(2)}$, we provide below a remarkable structural property for LENs.
@@
We note that this constraint is stricter than the one identified for GENs. A GEN requires the symmetry for boundary hyperplanes while LEN requires that for channel vectors, where one boundary hyperplane has to be determined by at least one channel vector but different channel vectors could leads to one hyperplane. For instance, the channel vectors $\alpha$ and $\alpha$ are distinct vectors but define the exact same boundary hyperplane. A GEN has the flexibility to use scaled versions of the same vector normal, whereas a LEN, in its symmetric form, must treat each parallel vector as a distinct channel vector.
@@ 5 Drawback of Equivariance in Expressive Power
In this section, we show our results on how enforcing equivariance hurts the expressive power by constructing a target function on $\mathbb{R}^{2}\to\mathbb{R}$. When approximating this target function, we show that for neural networks with one neuron, i.e. in parameter set $\Theta_{1}$, GNs have strictly smaller expected loss than GENs. For 3-neuron-networks, GNs and GENs can fully approximate this target function, while LENs can not.
@@
$s\circ\rho_{g}\begin{pmatrix}a\\ b\end{pmatrix}=s\begin{pmatrix}b\\ a\end{pmatrix}=\max(|b|,|a|,|b-a|)=s\begin{pmatrix}a\\ b\end{pmatrix}$,
@@
Figure 1 provides a 3D plot and a 2D contour map to visualize the effects of the imposed symmetry. Furthermore, Figure 2 displays the feature function of the symmetry group element sss. Building on this example, we demonstrate how the constraints of GENs can limit expressive power and how this limitation is further pronounced in LENs. Specifically, we mainly focus on GENs vs GNs on the 1-neuron setting and LENs vs GENs on the 3-neuron setting.
For one-neuron networks
the output function of the neural network becomes $\beta\sigma(\langle\alpha,x\rangle)$. We consider GNs first, where no equivariance is required on the network. One trivial one-neuron network is the ReLU output of every dimension of input. Formally, when $\alpha=(1,0)^{\top}$ and $\beta=1$, we have $F((a,b)^{\top})=\sigma(a)$ being an approximation and $F$ is a GN with one neuron. Similarly when $\alpha=(0,1)^{\top}$ and $\beta=1$, $F((a,b)^{\top})=\sigma(b)$ is also a GN approximation. Thus, for all input distribution $P$ with $\operatorname{Pr}_{x\sim P}[x^{\top}=(0,0)]<1$, we have
@@
For GENs, which has property $f(\rho x)=f(x)$ for $f(x)=\langle\mathcal{F}(x),x\rangle$ and $\rho$ being $\rho_{g}$ and $\rho_{h}$. Then we have $\langle\mathcal{F}(\rho x),\rho x\rangle=\langle\rho^{T}\mathcal{F}(\rho x),x\rangle=\langle\mathcal{F}(x),x\rangle$, which implies $\rho^{T}\circ\mathcal{F}\circ\rho=\mathcal{F}$. As a result, we have for all $x$:
@@
where at least one of $\sigma(\langle\alpha,x\rangle)$ and $\sigma(\langle\alpha,-x\rangle)$ is zero, indicating $\mathcal{F}=0$ for all $x$. It shows that the only one-neuron GEN is zero, which leads to strictly less expressive power then GNs. Specifically, we have for all input distribution $P$ with $\operatorname{Pr}_{x\sim P}[x^{\top}=(0,0)]<1$:
@@
For three-neurons networks
GNs can represent $s$ defined in Example 5.1 directly because the form $s((a,b)^{\top})=\sigma(a)+\sigma(-b)+\sigma(b-a)$ naturally fits the form of a three-neuron network. Due to the equivariance property of $s$, this specific network is also a GEN. Therefore, 3 neurons are sufficient for both GNs and GENs to fully approximate $s$.
However, for LENs, we show that 3 neurons are not enough; in fact, LENs require at least 6 neurons to fully approximate $s$. According to Theorem 4.6, for any LEN that approximates $s$, there exists a rewritten LEN with symmetric channel vectors of the same size that also perfectly approximates $s$. Due to the three boundary hyperplanes of $s$, this rewritten LEN must define three hyperplanes: $x=0$, $y=0$, and xy=0x-y=0, which implies the existence of at least three channel vectors: $c_{1}(1,0)$, $c_{2}(0,1)$, and $c_{3}(1,1)$ with $c_{1}c_{2}c_{3}\neq 0$. Since the channel vectors are symmetric, we have $\rho_{h}^{\top}c_{1}(1,0)=-c_{1}(1,0)$ is also a channel vector, so do $c_{2}(0,1)$ and $c_{3}(1,1)$. Consequently, the origin LEN must also contain at least 6 channel vectors, which applies to any LEN that fully approximates $s$.
Finally, we confirm that 6 neurons are sufficient, as the following LEN fully approximates $s$:
$F_{\theta}((a,b)^{\top})=\frac{1}{2}[\sigma(a)+\sigma(-a)+\sigma(b)+\sigma(-b)+\sigma(b-a)+\sigma(a-b)]$
@@ 6 Compensation to Enforcing Equivariance
In the previous sections, we demonstrated that for a fixed model size, the expressive power of LENs may strictly less than that of GENs, which are also strictly less expressive than GNs. A natural question arises: can this limitation in expressive power be overcome by increasing the model size? This section affirms that increasing model size is indeed a viable solution. We will examine this through both realizable and non-realizable cases. Furthermore, we also show that a key advantage of equivariant networks is their reduced hypothesis complexity, which leads to better generalizability.
6.1 Realizable case: Enlarging model size of LENs to compensate expressive power
We first consider the realizable case, where the target function can be perfectly represented by a GN with a finite number of neurons. By definition, this also implies that $s$ can be represented by a finite-neuron GEN. Our objective is to demonstrate the ability of LEN to approximate such a finite-neuron GEN.
Firstly, we show that with a larger model size, there always exists a LEN that can represent any representable mapping of finite-neuron GENs. To demonstrate this, consider a target GEN function defined as $s(x)=\sum_{i=1}^{m}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$. We can construct a LEN, denoted by $F(x)$, that perfectly replicates $s(x)$. The weights for this constructed LEN are defined as follows:
$W^{(2)}=\frac{1}{|G|}[\underbrace{\beta_{1},\dots,\beta_{1}}_{|G|},\dots,\underbrace{\beta_{m},\dots,\beta_{m}}_{|G|}]~~~~~\text{and}$
@@ 6.1 Realizable case: Enlarging model size of LENs to compensate expressive power
This construction ensures that the weights adhere to the necessary equivariance conditions. For the first layer, we have for all $g\in G$, $\rho_{g_{i}}\left\langle\alpha_{k},\rho_{g_{i}}(x)\right\rangle=\left\langle\rho_{g_{i}}^{\top}\rho_{g_{i}}^{\top}\alpha_{k},x\right\rangle=\left\langle\rho_{g_{i}g_{j}}^{\top}\alpha_{k},x\right\rangle$, which implies the existence of a permutation $\psi$ that maps the neuron index corresponding to $\rho_{g_{i}}^{\top}\alpha_{k}$ to the index for $\rho_{g_{i}g_{j}}^{\top}\alpha_{k}$. This $\psi$ is admitted because of Theorem 4.4. For $W^{(2)}$, we have the $\beta$ for all $\rho_{h}^{\top}\alpha_{i}$ is always $\beta$, indicating its invariance with respect to $\psi$.
This constructed LEN is a rewritten version of $s$ because of its invariance and equivariance. Specifically, we have
@@
which confirms that our constructed LEN perfectly represents the target function $s$. This principle can also be extended to handle equivariant functions and multi-layer architectures, as detailed in Appendix E.
Building on this, we utilize the relationship between the number of boundary hyperplanes and that of channel vectors to study how many neurons are required to let LEN approach GEN. For GENs/LENs, the number of channel vectors are larger than the number of boundary hyperplanes by Theorem 4.2. Inversely for GENs, if there are $k$ boundary hyperplanes of some output function $k$, we can make some specific $\beta_{i}=1$ and $\alpha_{i}$ for $i\in[k]$ such that $f-\sum_{i=1}^{k}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$ has no boundary hyperplanes, which is then can be represented as a linear function $\langle\alpha_{0},x\rangle=\sigma(\langle\alpha_{0},x\rangle)-\sigma(\langle-\alpha_{0},x\rangle)$. Therefore, the number of boundary hyperplanes is larger than the number of channel vectors plus two for GENs. As for LENs, we utilize the following lemma to deal with all symmetric channel vectors as a whole part.
@@
Let $\sum_{i=1}^{\ell}\beta_{i}\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\alpha_{i},x\rangle)$ be an output function of LENs with $\alpha_{i}\not=0$ for all $i\in[\ell]$. Denote $M_{i}=\{\rho_{g}^{T}\alpha_{i}/\|\rho_{g}^{T}\alpha_{i}\|:g\in G\}$ as the normalized orbit induced by $\alpha_{i}$ and then we have $M_{i}=M_{j}$ or $M_{i}\cap M_{j}=\phi$ for all $i,j\in[\ell]$. Moreover, if $M_{i}=M_{j}$, we have
@@
which enables us to reformulate the output function as $\sum_{i=1}^{\ell^{\prime}}\tilde{\beta_{i}}\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\tilde{\alpha_{i}},x\rangle)$ such that $\tilde{M}_{i}\cap\tilde{M}_{j}=\phi$ for all $i,j\in[\ell^{\prime}]$. Furthermore, for each $\tilde{M}_{i}$, we can then merge all parallel channel vectors since
@@
6.2 Non-realizable case: Enlarging model size can lead to even better expressive power
We now consider the non-realizable problem. With the same model size, equivariance may hurt the expressive power. We prove that when the networks have a larger model size, LENs may achieve a comparable and even better expressive power. We start with the following theorem, which shows a comparable expressive power when the model size is enlarged $|G|$ times.

### v2

@@ 4.2 LENs further Imply Symmetric Channel Vectors
We first prove that each row of $\psi_{g}$ has at most one non-zero entry for all $g\in G$. Assume for contradiction that exists $g\in G$ where a row $i$ of $\psi_{g}$ has at least two non-zero entries, $(\psi_{g})_{ij}$ and $(\psi_{g})_{ik}$. Without loss of generality, we assume $(\psi_{g})_{ij}\leq(\psi_{g})_{ik}$. Admitted $\psi$ means that for input vector $x=e_{j}-e_{k}$, we have $(\sigma(\psi_{g}x))_{i}=(\psi_{g}\sigma(x))_{i}=(\psi_{g}e_{j})_{i}=(\psi_{g})_{ij}$. However, on the other side we have $(\sigma(\psi_{g}x))_{i}=\sigma({(\psi_{g})_{ij}-(\psi_{g})_{ik}})=0$, which contradicts our initial assumption that the entries were non-zero. Therefore, each row of $\psi_{g}$ contains at most one non-zero entry.
@@
Consider a pair $(i,j)$ that makes $\psi_{ij}\neq 0$ for some $g\neq h$. Admitted $\psi$ means that for input $e_{j}$, we have $\beta_{j}=W^{(2)}e_{j}=W^{(2)}\psi e_{j}=\psi_{ij}W^{(2)}e_{i}=\psi_{ij}\beta_{i}$ for all $\psi$, including $\psi_{g}$ and $\psi_{h}$. It shows that $\beta_{i}$ and $\beta_{j}$ share the same direction, meaning $\|\beta_{j}\|\beta_{i}=\|\beta_{i}\|\beta_{j}$. It also shows that $\psi_{ij}$ does not change if $\psi$ switches from $\psi_{g}$ to $\psi_{h}$, i.e., $(\psi_{g})_{ij}=(\psi_{h})_{ij}$. ∎
In light of Lemma 4.5, we say $\beta_{i}$ and $\beta_{j}$ are in the same orbit if they share the same direction, i.e., $\|\beta_{j}\|\beta_{i}=\|\beta_{i}\|\beta_{j}$. The lemma demonstrates that the symmetry constraint requires all weight vectors in the same orbit to lie in the same direction in vector space.
Aside from the constraint on $W^{(2)}$, LENs also impose a condition on $W^{(1)}$ - $W^{(1)}$ should belong to the intertwiner space defined by $\psi$ and $\rho$. Combining it with the constraints on $\psi$ and $W^{(2)}$, we provide the following theorem.
@@
We note that this constraint is stricter than the one for GENs. A GEN requires the symmetry for boundary hyperplanes while LEN requires that for channel vectors, where one boundary hyperplane has to be determined by at least one channel vector, but different channel vectors could lead to one hyperplane. For instance, the channel vectors $\alpha$ and $\alpha$ are distinct vectors but define the exact same boundary hyperplane. A GEN has the flexibility to use scaled versions of the same vector normal, whereas a LEN, in its symmetric form, must treat each parallel vector as a distinct channel vector.
@@ 5 Drawback of Equivariance in Expressive Power
In this section, we show our results on how enforcing equivariance hurts the expressive power. The proofs are through constructing a target function on $\mathbb{R}^{2}\to\mathbb{R}$. When approximating this target function, we show that for neural networks with one neuron, i.e. in parameter set $\Theta_{1}$, GNs have strictly smaller expected loss than GENs. For 3-neuron-networks, GNs and GENs can fully approximate this target function, while LENs cannot.
@@
$s\circ\rho_{g}\begin{pmatrix}a\\ b\end{pmatrix}=s\begin{pmatrix}b\\ a\end{pmatrix}=\max(|b|,|a|,|b-a|)=s\begin{pmatrix}a\\ b\end{pmatrix},\text{ and}$
@@
Figure 1 provides a 3D plot and a 2D contour map to visualize the effects of the imposed symmetry. Figure 2 displays the feature function of the symmetry group element. Building on this example, we demonstrate how the constraints of GENs can limit expressive power and how this limitation is further pronounced in LENs. Specifically, we mainly focus on comparisons of GENs vs. GNs on the 1-neuron setting, and LENs vs GENs on the 3-neuron setting.
One-neuron networks.
The output function of the neural network becomes $\beta\sigma(\langle\alpha,x\rangle)$. A typical one-neuron network applies the ReLU activation function to each dimension of the input. Formally, when $\alpha=(1,0)^{\top}$ and $\beta=1$, $F((a,b)^{\top})=\sigma(a)$ is an approximation and $F$ is a GN with one neuron. Similarly, when $\alpha=(0,1)^{\top}$ and $\beta=1$, $F((a,b)^{\top})=\sigma(b)$ is also a GN approximation. Thus, for all input distribution $P$ with $\operatorname{Pr}_{x\sim P}[x^{\top}=(0,0)]<1$, we have
@@
For GENs, we have $f(\rho x)=f(x)$. Combining with the definition of feature function $\mathcal{F}$, we have $\langle\mathcal{F}(\rho x),\rho x\rangle=\langle\rho^{T}\mathcal{F}(\rho x),x\rangle=\langle\mathcal{F}(x),x\rangle$, which implies $\rho^{T}\circ\mathcal{F}\circ\rho=\mathcal{F}$ for all $\rho$. Let $\rho=\rho_{h}$, we have for all $x$:
@@
where at least one of $\sigma(\langle\alpha,x\rangle)$ and $\sigma(\langle\alpha,-x\rangle)$ is zero, indicating $\mathcal{F}=0$ for all $x$. It shows that the only one-neuron GEN is zero, which leads to strictly lower expressive power then GNs. Specifically, we have that for all input distribution $P$ with $\operatorname{Pr}_{x\sim P}[x^{\top}=(0,0)]<1$,
@@
Three-neuron networks.
Three-neuron GNs can represent $s$ defined in Example 5.1 because the form $s((a,b)^{\top})=\sigma(a)+\sigma(-b)+\sigma(b-a)$ is already a three-neuron network. Due to the equivariance property of $s$, this network is also a GEN. Therefore, three neurons are sufficient for both GNs and GENs to achieve $s$.
However, for LENs, we show that three neurons are not enough. According to Theorem 4.6, for any LEN that represents $s$, there exists a rewritten LEN with symmetric channel vectors of the same size that also perfectly represents $s$. Since $s$ has three boundary hyperplanes: $x=0$, $y=0$, and xy=0x-y=0, they should also be the boundary hyperplane of the rewritten LEN, which implies the existence of at least three channel vectors: $c_{1}(1,0)$, $c_{2}(0,1)$, and $c_{3}(1,1)$ with $c_{1}c_{2}c_{3}\neq 0$. Since the channel vectors are symmetric, $\rho_{h}^{\top}c_{1}(1,0)=-c_{1}(1,0)$also forms a channel vector; as do $c_{2}(0,1)$ and $c_{3}(1,1)$. As a result, the original LEN must contain at least six channel vectors, i.e., LENs require at least six neurons to represent $s$. Finally, we show that six neurons are sufficient, as the following LEN has exact six neurons and represents $s$:
$F_{\theta}((a,b)^{\top})=\frac{1}{2}[\sigma(a)+\sigma(-a)+\sigma(b)+\sigma(-b)+\sigma(b-a)+\sigma(a-b)].$
@@ 6 Compensation to Enforcing Equivariance
In the previous sections, we demonstrated that for a fixed model size, the expressive power of LENs are strictly lower than that of GENs, which are also strictly less expressive than GNs. A natural question arises: can this limitation in expressive power be compensated for by enlarging the model size? This section affirms that enlarging the model size is indeed a viable solution. We will examine this through both realizable and non-realizable cases. Further, we show that, surprisingly, the enlarged networks can have even reduced hypothesis complexity, which leads to better generalizability.
6.1 Realizable Case
We first consider the realizable case, where the target function can be perfectly represented by a GN with a finite number of neurons. By definition, $s$ can also be represented by a finite-neuron GEN. When it comes to LEN, the conclusion is not straightforward.
Firstly, we show that with a larger model size, there always exists a LEN that can represent any representable mapping of finite-neuron GENs. To demonstrate this, consider a target GEN-representable function defined as $s(x)=\sum_{i=1}^{m}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$. We can construct a LEN, denoted by $F(x)$, that perfectly replicates $s(x)$. The weights for this constructed LEN are defined as follows:
$W^{(2)}=\frac{1}{|G|}[\underbrace{\beta_{1},\dots,\beta_{1}}_{|G|},\dots,\underbrace{\beta_{m},\dots,\beta_{m}}_{|G|}],\text{ and}$
@@ 6.1 Realizable Case
This construction ensures that the weights adhere to the necessary equivariance conditions. In the first layer, for all $g\in G$, $\rho_{g_{i}}\left\langle\alpha_{k},\rho_{g_{i}}(x)\right\rangle=\left\langle\rho_{g_{i}}^{\top}\rho_{g_{i}}^{\top}\alpha_{k},x\right\rangle=\left\langle\rho_{g_{i}g_{j}}^{\top}\alpha_{k},x\right\rangle$, implying the existence of a permutation $\psi$ that maps the neuron index corresponding to $\rho_{g_{i}}^{\top}\alpha_{k}$ to the index for $\rho_{g_{i}g_{j}}^{\top}\alpha_{k}$. This $\psi$ is admitted because of Theorem 4.4. In the second layer, we have that the $\beta$ is $\beta$ for all $\rho_{h}^{\top}\alpha_{i}$, indicating its invariance with respect to $\psi$.
This constructed LEN is a rewritten version of $s$ because of its invariance and equivariance; specifically,
@@
which confirms that our constructed LEN perfectly represents the target function $s$. This result can be extended to handle any equivariant functions and multi-layer architectures, as detailed in Appendix E.
Building on this, we utilize the relationship between the number of boundary hyperplanes and that of channel vectors to study how many neurons are required to let LEN approach GEN. For GENs or LENs, the number of channel vectors is larger than the number of boundary hyperplanes, according to Theorem 4.2. Inversely, for $f\in\operatorname{GENs}$, if there are $k$ boundary hyperplanes of some output function $k$, we can construct some specific $\beta_{i}=1$ and $\alpha_{i}$ for $i\in[k]$ such that $f-\sum_{i=1}^{k}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$ has no boundary hyperplanes, which then can be represented as a linear function $\langle\alpha_{0},x\rangle=\sigma(\langle\alpha_{0},x\rangle)-\sigma(\langle-\alpha_{0},x\rangle)$. Therefore, for GENs, we can compress neurons to make the number of channel vectors equal to the number of boundary hyperplanes plus two.
As for LENs, we prove the following lemma that deals with all symmetric channel vectors as a whole.
@@
Let $\sum_{i=1}^{\ell}\beta_{i}\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\alpha_{i},x\rangle)$ be an output function of LENs with $\alpha_{i}\not=0$ for all $i\in[\ell]$. Denote $M_{i}=\{\rho_{g}^{T}\alpha_{i}/\|\rho_{g}^{T}\alpha_{i}\|:g\in G\}$ as the normalized orbit induced by $\alpha_{i}$. Then, we have $M_{i}=M_{j}$ or $M_{i}\cap M_{j}=\phi$ for all $i,j\in[\ell]$. Moreover, if $M_{i}=M_{j}$, we have
@@
which enables us to reformulate the output function as $\sum_{i=1}^{\ell^{\prime}}\tilde{\beta_{i}}\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\tilde{\alpha_{i}},x\rangle)$ such that $\tilde{M}_{i}\cap\tilde{M}_{j}=\phi$ for all $i,j\in[\ell^{\prime}]$. Furthermore, for each $\tilde{M}_{i}$, we can then merge all parallel channel vectors, since
@@
6.2 Non-realizable Case
We now consider the non-realizable problem. With the same model size, enforcing equivariance may hurt the expressive power. We prove that when the networks have a larger model size, LENs may achieve a comparable and even better expressive power. We start with the following theorem, which shows a comparable expressive power when the model size is enlarged $|G|$ times.